# Key-Point-Driven Mathematical Reasoning Distillation of Large Language Model

### v1

@@ Abstract
Xunyu Zhu1,2, Jian Li1,2thanks: Corresponding author, Yong Liu3, Can Ma1,2, Weiping Wang1,2 1Institute of Information Engineering, Chinese Academy of Sciences 2School of Cyber Security, University of Chinese Academy of Sciences 3Gaoling School of Artificial Intelligence, Renmin University of China {zhuxunyu, lijian9026, macan, wangweiping}@iie.ac.cn, liuyonggsai@ruc.edu.cn
@@ 1 Introduction
Large language models (LLMs) based on Transformer architectures represent a significant advancement in natural language processing. Notable models such as LLaMA (Touvron et al., 2023a), GPT-4 (OpenAI, 2023), and PaLM (Chowdhery et al., 2023) feature hundreds of billions of parameters. Trained on extensive text datasets, these models exhibit exceptional proficiency across a broad range of downstream tasks.
Recent studies (Chen et al., 2023; Wang et al., 2023b, a; Liu et al., 2023) have enhanced the mathematical reasoning abilities of LLMs through Chain-of-Thought (CoT) prompting, which generates intermediate steps to solve complex problems. However, deploying these models remains challenging due to their size and computational demands. For instance, the GPT-3 model (Brown et al., 2020) requires at least 350GB of FP16 storage and multiple A100 GPUs with 80GB of memory each for efficient inference.
@@
Recent work (Magister et al., 2023; Shridhar et al., 2023; Ho et al., 2023; Fu et al., 2023) investigates distilling LLM reasoning into SLMs (under 1B parameters) for broader deployment. This involves using LLMs to create enriched datasets with detailed reasoning paths, which then fine-tune SLMs, endowing them with advanced reasoning abilities. For example, Chain-of-Thought Distillation (CoTD)(Ho et al., 2023) encapsulates the reasoning process into textual rationales. However, there is a significant performance gap between SLMs and LLMs. Prior work(Wei et al., 2022) identifies three main error types in CoT reasoning: (1) Calculation errors: Incorrect calculations leading to wrong answers. (2) Missing Step errors: Omissions of intermediate reasoning steps, especially in multi-step problems. (3) Semantic misunderstanding errors: Errors in understanding the problem or maintaining coherent reasoning, often due to insufficient model capability. To explore the reasons for the performance gap between SLMs and LLMs, we conducted the same error analysis on CoTD. Our preliminary experiments (shown in Figure 1) reveal numerous error combinations in CoTD, with calculation and semantic misunderstanding errors being the most prevalent. Prior work Zhu et al. (2023a) proposed Program-of-Thought Distillation (PoTD) to mitigate calculation errors by formulating the reasoning process as a Python program executed by an external interpreter. This approach allows the SLM to focus on generating the program, avoiding calculation errors and improving reasoning performance. Given these circumstances, our paper focuses on addressing semantic misunderstanding errors in CoTD to further enhance the reasoning performance of SLMs.
@@
We assessed KPDD across FlanT5 models from Small (0.06B) to Large (0.76B) on four mathematical reasoning datasets. The results show that KPDD-CoT significantly enhances SLMs’ reasoning abilities, while KPDD-PoT enables SLMs to achieve state-of-the-art (SOTA) mathematical reasoning performance. For instance, with KPDD-CoT, FlanT5-Large achieved an average accuracy of 24.71% on these datasets, and KPDD-PoT elevated FlanT5-Large to an average accuracy of 63.83%. Furthermore, our error analysis on KPDD confirms that KPDD effectively mitigates misunderstanding errors, thereby improving the mathematical reasoning performance of SLMs.
@@ 2 Related Work
2.1 Mathematical Reasoning
Mathematical reasoning tasks, exemplified by benchmarks such as GSM8K (Cobbe et al., 2021) and SVAMP (Patel et al., 2021), present a substantial challenge for LLMs. To enhance LLMs’ performance in this domain, researchers have identified two primary strategies.
Chain-of-Thought Reasoning LLMs’ reasoning ability can be enhanced by prompting them to articulate intermediate steps towards a solution, as demonstrated by Wei et al. Wei et al. (2022). This insight has spurred various advancements (Chen et al., 2023; Wang et al., 2023b, a; Liu et al., 2023) that refine reasoning paths: Chen et al. Chen et al. (2023) prompt LLMs to generate executable code; Wang et al. Wang et al. (2023b) use multiple reasoning paths with a voting mechanism; Wang et al. Wang et al. (2023a) have LLMs create a plan before reasoning; Liu et al. Liu et al. (2023) employ diverse reasoning prompts for problem-solving; Zhong et al. Zhong et al. (2024) encourage LLMs to deeply understand problems and leverage key information for better reasoning. Building on these methods, our work introduces Key-Point-Driven Mathematical Reasoning Distillation (KPDD) to further enhance SLMs’ mathematical reasoning.
Finetuning-based Reasoning refines LLMs such as Llama2 (Touvron et al., 2023b), Qwen (Bai et al., 2023), and Baichuan2 (Yang et al., 2023) by integrating techniques from advanced models like GPT-4 (OpenAI, 2023) and PaLM-2 (Anil et al., 2023). Notably, Yuan et al.Yuan et al. (2023) utilize Rejection Sampling Fine-Tuning (RFT) to enhance LLMs’ mathematical reasoning, while WizardMath(Luo et al., 2023) employs Reinforcement Learning from Evolved Instructions Feedback (RLEIF) to improve LLaMA-2’s reasoning abilities. MAmmoTH (Yue et al., 2023) combines Chain-of-Thought (CoT) and Program-of-Thought (PoT) rationales for more effective instruction-tuning of LLMs in math problem-solving. Despite their effectiveness, the large model sizes of these LLMs continue to limit their deployment efficiency.
2.2 Knowledge Distillation
Knowledge Distillation optimizes LLMs for practical use by transferring knowledge from larger models to smaller, efficient ones (Zhu et al., 2023b). Recent research (Magister et al., 2023; Shridhar et al., 2023; Ho et al., 2023; Fu et al., 2023) focuses on endowing compact models ($\leq$ 1B parameters) like T5 (Raffel et al., 2020) and GPT-2 (Radford et al., 2019) with advanced reasoning capabilities from LLMs such as GPT-4 (OpenAI, 2023) and PaLM-2 (Anil et al., 2023). For instance, Ho et al. Ho et al. (2023) fine-tune student models using accurate reasoning paths from LLMs, Shridhar et al. Shridhar et al. (2023) train dual-model systems on sub-questions and solutions, and Fu et al. Fu et al. (2023) propose scaling down general competencies of smaller models to enhance task-specific performance. Our work introduces a novel distillation approach where two SLMs independently extract the core question and key problem-solving information from an original question. These key points are then utilized to guide another SLM in solving the original question effectively.
@@ 3 Method
Figure 2: Detailed data generation for KPDD-CoT. Firstly, we manually construct some contextualized examples, and combine these contextualized examples, the question, and the prompt "Firstly, let’s extract the most comprehensive and detailed key question. Then, let’s identify and list the most useful information related to the question. Finally, let’s understand the key question and the problem-solving information, solve the question step by step, and show the answer." to prompt LLMs to generate the corresponding reasoning process. Then, we extract the answer from this reasoning process. If the answer doesn’t agree with the gold answer, we will drop the reasoning process. Finally, we get a high-quality KPDD-CoT reasoning dataset.
@@ 3.1 Data Generation from LLMs
Our KPDD method begins by creating a mathematical reasoning dataset from LLMs, which is then used for SLM fine-tuning. In our paper, we use in-context learning (Dong et al., 2023; Min et al., 2022; Rubin et al., 2022) to prompt LLMs for constructing the reasoning dataset. Furthermore, in stage 3, our KPDD method employs two distillation approaches: one distills the SLM to generate CoT rationales for problem-solving, and the other distills the SLM to generate PoT rationales for problem-solving. In other words, our KPDD method can be divided into two approaches: KPDD-CoT and KPDD-PoT.
@@ 3.1.1 Data Generation for KPDD-CoT
Given a mathematical dataset $\mathcal{D}$, each entry $(x,y)$ pairs a question $x$ with its answer $y$. As illustrated in Figure 2, we select $k$ samples $\{(x_{1},y_{1}),\ldots,(x_{k},y_{k})\}$ from $\mathcal{D}$ and manually craft reasoning processes. Each reasoning process $c$ includes a core question, problem-solving information, and rationales in CoT format. These elements are separated by HTML tags: "<core>{core question}</core><info>{problem-solving information}</info><cot>{rationales in CoT format}</cot>". These form contextualized instances $\{(x_{1},c_{1},y_{1}),\ldots,(x_{k},c_{k},y_{k})\}$, compiled into a demonstration set $\mathcal{D}_{c}$. We then prompt LLMs with the demonstration set $\mathcal{D}_{c}$, a new question, and the instruction "Firstly, let’s extract the most comprehensive and detailed key question. Then, let’s identify and list the most useful information related to the question. Finally, let’s understand the key question and the problem-solving information, solve the question step by step, and show the answer." to generate the reasoning process for the new question. The KPDD-CoT dataset generation is formalized as:
@@
where $\mathcal{M}$ denotes the LLM, $f$ is the decoding function, and $i$ denotes the index in $\mathcal{D}$. This yields the KPDD-CoT dataset $\mathcal{D}_{C}$, composed of triplets $(x,c,y)$.
Data Filtering—Upon generating the KPDD-CoT dataset with LLMs, we validate the reasoning process against the gold standard answer—a crucial step to ensure the quality of our reasoning dataset $\mathcal{D}_{C}$. Discrepancies between the generated reasoning process and the gold standard answer result in the exclusion of those entries from $\mathcal{D}_{C}$. This meticulous filtering removes incorrect examples, thereby enhancing the dataset’s overall quality. Finally, this refinement directly contributes to the improved performance of fine-tuned SLMs, due to the increased accuracy and reliability of the training data. By ensuring that only high-quality reasoning processes are included, we bolster the effectiveness of the SLMs in solving mathematical reasoning tasks.
Figure 3: Detailed data generation for KPDD-PoT. Firstly, we manually construct some contextualized examples, and combine these contextualized examples, the question, and the prompt "Firstly, let’s extract the most comprehensive and detailed key question. Then, let’s identify and list the most useful information related to the question. Finally, let’s understand the key question and the problem-solving information, and generate the python code (return ans) to solve the question." to prompt LLMs to generate the corresponding reasoning process. Then, we extract the rationale in PoT format from this reasoning process, and send the rationale to a extra python interpreter. If there are compile errors or if it produces wrong answer, we will drop the reasoning process. Finally, we get a high-quality KPDD-PoT reasoning dataset.
@@ 3.1.2 Data Generation for KPDD-PoT
Similar with KPDD-CoT, the initial phase in our KPDD-PoT entails creating a dataset from LLMs, setting the stage for SLM fine-tuning. For KPDD-CoT dataset generation, we also choose $k$ samples $\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{k},y_{k})\}$ from $\mathcal{D}$ and manually create reasoning processes $p$, where each reasoning process includes a core question, problem-solving information, and rationales in PoT format. These elements are also separated by HTML tags: "<core>{core question}</core><info>{problem-solving information}</info><pot>{rationales in PoT format}</pot>". These form contextualized instances $\{(x_{1},p_{1},y_{1}),(x_{2},p_{2},y_{2}),\ldots,(x_{k},p_{k},y_{k})\}$, which are compiled into a demonstration set $\mathcal{D}_{p}$. We then prompt the LLM with the demonstration set $\mathcal{D}_{p}$ and a new question, and input the instruction "Firstly, let’s extract the most comprehensive and detailed key question. Then, let’s identify and list the most useful information related to the question. Finally, let’s understand the key question and the problem-solving information, and generate the python code (return ans) to solve the question." to generate a reasoning process for the new question. Figure 3 outlines this data generation process, and the KPDD-PoT dataset generation is formalized as:
$p_{i}=f_{\mathcal{M}}(x_{i},\mathcal{D}_{p}),$ (2)
where $\mathcal{M}$ is the LLM, $f$ denotes the greedy decoding function, and $i$ is represented as the index of the instance $(x,y)$ in $\mathcal{D}$. This yields a KPDD-PoT dataset $\mathcal{D}_{P}$, organized as triplets $(x,p,y)$.
Data Filtering—Following KPDD-PoT dataset generation by LLMs, each program in the reasoning process undergoes validation using an external Python interpreter, a vital step to ensure the quality of our initial dataset $\mathcal{D}_{P}$. Programs that fail to compile or produce incorrect results are immediately discarded. This rigorous filtering process removes flawed instances, thus improving the dataset’s quality.
@@ 3.2.1 Fine-tuning SLMs for KPDD-CoT
Then, we construct a problem-solving subset from the KPDD-CoT dataset, denoted as $\mathcal{D}_{CI}$. Each sample in this subset can be represented as $(x,ci)$, where $x$ represents the original question and $ci$ represents the problem-solving information. For each training instance $(x,ci)$ from $\mathcal{D}_{CI}$, we prepend the prompt $p_{ci}$Let’s identify and list the most useful information related to the question." to the question $x$. This guides the KPDD-CoT-info in fine-tuning to accurately extract the corresponding problem-solving information $ci$. The fine-tuning loss function can be represented as follows:
@@
Finally, we construct a problem-solving subset from the KPDD-CoT dataset, denoted as $\mathcal{D}_{CS}$. Each sample in this subset can be represented as $(x,cc,ci,cs)$, where $x$ represents the original question, $cc$ represents the core question, $ci$ represents the problem-solving information, and $cs$ represents the rationales in CoT format. For each training instance $(x,cc,ci,cs)$ from $\mathcal{D}_{CS}$, we integrate the original question $x$, the core question $cc$, the problem-solving information $ci$ and the prompt $p_{cs}$Let’s understand the core question and the problem-solving information, solve the question step by step, and show the answer." to construct a new input. This guides the KPDD-CoT-solve in fine-tuning to generate rationales $cs$ for solving the origin question in CoT format. The fine-tuning loss function can be represented as follows:
@@ 4.1 Dataset
In our paper, the training dataset is derived from the GSM8K training set, which comprises diverse grade school math word problems (Cobbe et al., 2021). Additionally, the mathematical reasoning capabilities of the SLMs are evaluated using the GSM8K test set, along with other datasets including ASDiv, which contains diverse math word problems (Miao et al., 2020), SVAMP, which features math word problems with varying structures (Patel et al., 2021), and MultiArith, which consists of arithmetic word problems (Roy and Roth, 2015). The statistics of these datasets are summarized in Table 1. This comprehensive evaluation approach ensures that the SLMs’ mathematical reasoning capabilities are thoroughly tested across a variety of problem types and structures, providing a robust assessment of their performance.

### v2

@@ Abstract
Xunyu Zhu1,2, Jian Li1,2thanks: Corresponding author, Can Ma1,2, Weiping Wang1,2 1Institute of Information Engineering, Chinese Academy of Sciences 2School of Cyber Security, University of Chinese Academy of Sciences {zhuxunyu, lijian9026, macan, wangweiping}@iie.ac.cn
@@ 1 Introduction
Large language models (LLMs) have achieved impressive performance in mathematical reasoning tasks. Recent work (Wang et al., 2023) further proposes Chain-of-Thought (CoT) to enhance the mathematical reasoning abilities of LLMs. However, the massive scale of LLMs presents significant challenges for deployment.
@@
A feasible solution to address this problem is to use black-box distillation to transfer mathematical reasoning abilities from LLMs to SLMs. Chain-of-Thought Distillation (CoTD) Ho et al. (2023) is a representative mathematical reasoning distillation method. CoTD prompts LLMs to generate reasoning processes for each question, constructing a mathematical reasoning dataset. This dataset is then used to fine-tune SLMs, enhancing their mathematical reasoning abilities. However, there remains a significant performance gap between SLMs and LLMs. Prior work (Wei et al., 2022) identifies three main error types in CoT reasoning: Calculation errors, Missing Step errors, and Semantic misunderstanding errors. To explore the reasons for the performance gap between SLMs and LLMs, we conducted the same error analysis on CoTD. Our preliminary experiments (shown in Figure 1) reveal numerous error combinations in CoTD, with calculation and semantic misunderstanding errors being the most prevalent. Prior work Zhu et al. (2023) proposed Program-of-Thought Distillation (PoTD) to avoid calculation errors by formulating the reasoning process as a Python program executed by an external interpreter. This approach allows the SLM to focus on generating the program, avoiding calculation errors and improving reasoning performance. Given these circumstances, our paper focuses on addressing semantic misunderstanding errors in CoTD to further enhance the reasoning performance of SLMs.
@@
We use KPDD to fine-tune FlanT5 models, and evaluate these models on several mathematical reasoning dataset, including GSM8K, ASDiv, SVAMP, and MultiArith. Our experiment results show that KPDD-CoT significantly enhances SLMs’ reasoning abilities, while KPDD-PoT enables SLMs to achieve state-of-the-art (SOTA) mathematical reasoning performance. Furthermore, our error analysis on KPDD confirms that KPDD effectively mitigates misunderstanding errors, thereby improving the mathematical reasoning performance of SLMs.
@@ 2 Related Work
2.1 Chain-of-Thought Reasoning
Chain-of-Thought refers to prompt LLMs to solve the question step by step. Prior work Wei et al. (2022) find that Chain-of-Thought can effectively improve the reasoning performance of LLMs. Based on the findings, Kojima et al. Kojima et al. (2022) further introduce zero-shot CoT, which significantly improves the model’s reasoning performance by simply adding the prompt "Let’s think step by step" before answering. To avoid calculation error in CoT, Chen et al. Chen et al. (2023) formulate the reasoning process into program. Furthermore, Wang et al. Wang et al. (2023) introduce a self-consistency decoding strategy, which generates diverse reasoning paths and then selects the most consistent answer by considering these paths comprehensively. Least-to-most prompting Zhou et al. (2023) breaks down complex problems into a series of simpler subproblems and solves these subproblems sequentially. Zhong et al. Zhong et al. (2024) encourage LLMs to deeply understand problems and leverage key information for better reasoning. Inspired by these methods, our work introduces Key-Point-Driven Mathematical Reasoning Distillation (KPDD) to further enhance SLMs’ mathematical reasoning.
2.2 Black-box Distillation
Currently, close-source LLMs usually have stronger reasoning performance than open-source LLMs. However, in general, we can only obtain the output of close-source LLMs. Based on this situation, black-box distillation is proposed to distill abilities from LLMs to SLMs. Specifically, black-box distillation first prompts LLMs to generate a distillation dataset, and then this dataset is used to fine-tune SLMs to improve their reasoning performance. For example, Ho et al. Ho et al. (2023) prompt LLMs to generate a CoT reasoning dataset, which was then used to fine-tune an SLM, thereby enhancing its reasoning ability. Shridhar et al. Shridhar et al. (2023) utilize a LLM to train a problem decomposer and a subproblem solver separately. They used the decomposer to break down the original problem into multiple subproblems, and then employed the subproblem solver to solve each subproblem individually. By integrating the solutions to these subproblems, they derived the final answer. Fu et al. Fu et al. (2023) analyze CoT distillation and find that there is a trade-off between the specific abilities and the general abilities of SLMs. When CoT distillation is used to enhance the specific abilities of SLMs, their general abilities decrease correspondingly. Zhu et al. Zhu et al. (2023) construct a distillation dataset, where the reasoning format changed from CoT reasoning to PoT reasoning. Through this method, the SLM can delegate the calculation process to an additional Python interpreter to avoid calculation errors, thereby enhancing the SLM’s mathematical reasoning performance. Zhu et al. Zhu et al. (2024) further formalize the reasoning process into equations and find that a diverse range of reasoning formats can effectively enhance the mathematical reasoning performance of SLMs. Our work introduces a novel distillation approach where two SLMs independently extract the core question and key problem-solving information from an original question. These key points are then utilized to guide another SLM in solving the original question effectively.
@@ 3 Method
Figure 2: Detailed data generation for KPDD-CoT. Initially, we use few-shot prompting to guide LLMs in producing reasoning processes. We then extract the correct answers from these processes, discarding any reasoning process that does not align with the correct answers. In this way, we have constructed a high-quality KPDD-CoT reasoning distillation dataset.
@@ 3.1 Data Generation from LLMs
When given a mathematical dataset, our primary task is to utilize an LLM to generate a reasoning process for each mathematical problem in the dataset. By doing so, we augment the mathematical dataset to construct the distillation dataset. Furthermore, in stage 3, our KPDD method employs two distillation approaches: one distills the SLM to generate CoT rationales for problem-solving, and the other distills the SLM to generate PoT rationales for problem-solving. In other words, our KPDD method can be divided into two approaches: KPDD-CoT and KPDD-PoT.
@@ 3.1.1 Data Generation for KPDD-CoT
Our method uses few-shot prompting to prompt the LLM to synthesize the reasoning dataset. Figure 2 reveals the distillation dataset generation process. Specifically, we first randomly sample $k$ data pairs $(x,y)$ from a mathematical dataset $\mathcal{D}$, where $x$ represents the question and $y$ represents the answer. Then, for these sampled data, we manually construct the reasoning process $c$. Each reasoning process $c$ includes a core question, problem-solving information, and rationales in CoT format. These elements are separated by HTML tags: "<core>{core question}</core><info>{problem-solving information}</info><cot>{rationales in CoT format}</cot>". Finally, we obtain a demonstration dataset $\mathcal{D}_{c}$. At the same time, we further introduce an instruction "Firstly, let’s extract the most comprehensive and detailed key question. Then, let’s identify and list the most useful information related to the question. Finally, let’s understand the key question and the problem-solving information, solve the question step by step, and show the answer.". By leveraging the demonstration set and the instruction, we prompt the LLM to generate the reasoning process for the mathematical questions. The KPDD-CoT dataset generation is formalized as:
@@
where $\mathcal{M}$ denotes the LLM, $f$ represents the decoding function, and $i$ is the index in $\mathcal{D}$.
Data Filtering—To further improve the quality of the dataset, we attempt to filter out incorrect reasoning processes from the KPDD-CoT dataset. Specifically, for each data point in the KPDD-CoT dataset, we extract the answer from its reasoning process. If this answer matches the gold answer, we retain the data point; otherwise, we discard it as the reasoning process is deemed incorrect. By ensuring that the KPDD-CoT dataset contains only correct reasoning processes, we enhance the LLM’s mathematical reasoning performance.
Figure 3: Detailed data generation for KPDD-PoT. Similar to KPDD-CoT, we prompt LLMs to generate the reasoning process in the KPDD-PoT format. We extract the rationale in the PoT format from this process and run it through a Python interpreter. If there are errors or incorrect answers, we discard the reasoning process. Finally, we constructed a high-quality KPDD-PoT reasoning dataset.
@@ 3.1.2 Data Generation for KPDD-PoT
Similar to KPDD-CoT, we first sample $k$ examples from a mathematical dataset $\mathcal{D}$, then use these samples to manually create a demonstration set $\mathcal{D}_{p}$. Each reasoning process in the demonstration set $\mathcal{D}_{p}$ includes a core question, problem-solving information, and rationales in PoT format. These elements are also separated by HTML tags: "<core>{core question}</core><info>{problem-solving information}</info><pot>{rationales in PoT format}</pot>". Then, we further introduce the instruction a instruction "Firstly, let’s extract the most comprehensive and detailed key question. Then, let’s identify and list the most useful information related to the question. Finally, let’s understand the key question and the problem-solving information, and generate the python code (return ans) to solve the question.". By utilizing the demonstration set $\mathcal{D}_{p}$ and the instruction, we prompt LLMs to generate reasoning processes in for questions. Figure 3 shows the data synthesis process for KPDD-PoT, and the KPDD-PoT dataset generation process can be formalized as:
$p_{i}=f_{\mathcal{M}}(x_{i},\mathcal{D}_{p}).$ (2)
Data Filtering—Similar to data filtering for KPDD-CoT, we also apply filtering to the KPDD-PoT dataset. Specifically, for each data point in the KPDD-PoT dataset, we extract the rationale in PoT format from its reasoning process. An external Python interpreter is then used to run the rationale. If the obtained answer do not match the gold answer, it indicate that the reasoning process is incorrect. Consequently, we remove the data point with the flawed reasoning process from the KPDD-PoT dataset, thereby enhancing the dataset’s quality.
@@ 3.2.1 Fine-tuning SLMs for KPDD-CoT
Then, we construct a problem-solving subset from the KPDD-CoT dataset, denoted as $\mathcal{D}_{CI}$. Each sample in this subset can be represented as $(x,ci)$, where $x$ represents the original question and $ci$ represents the problem-solving information. For each data point $(x,ci)$ from $\mathcal{D}_{CI}$, we add the prompt $p_{ci}$Let’s identify and list the most useful information related to the question." to the question $x$. This guides the KPDD-CoT-info in fine-tuning to accurately extract the corresponding problem-solving information $ci$. The fine-tuning loss function can be represented as follows:
@@
Finally, we construct a problem-solving subset from the KPDD-CoT dataset, denoted as $\mathcal{D}_{CS}$. Each sample in this subset can be represented as $(x,cc,ci,cs)$, where $x$ represents the original question, $cc$ represents the core question, $ci$ represents the problem-solving information, and $cs$ represents the rationales in CoT format. For each data $(x,cc,ci,cs)$ from $\mathcal{D}_{CS}$, we integrate the original question $x$, the core question $cc$, the problem-solving information $ci$ and the prompt $p_{cs}$Let’s understand the core question and the problem-solving information, solve the question step by step, and show the answer." to construct a new input. This guides the KPDD-CoT-solve in fine-tuning to generate rationales $cs$ for solving the origin question in CoT format. The fine-tuning loss function can be represented as follows:
@@ 4.1 Dataset
In our paper, we generate KPDD distillation datasets based on the GSM8K training set, which comprises diverse grade school math word problems (Cobbe et al., 2021). Then, we evaluate the mathematical reasoning performance of SLM on the GSM8K test set. Furthermore, to assess the transferability of SLMs’ mathematical reasoning capabilities, we evaluate SLM on several additional mathematical datasets. These datasets include ASDiv, which contains diverse math word problems (Miao et al., 2020), SVAMP, which features math word problems with varying structures (Patel et al., 2021), and MultiArith, which consists of arithmetic word problems (Roy and Roth, 2015). The statistics of these datasets are summarized in Table 1. This comprehensive evaluation approach ensures that the SLMs’ mathematical reasoning capabilities are thoroughly tested across a variety of problem types and structures, providing a robust assessment of their performance.